OpenStudy (anonymous):

If f(x)=(x+1)^-1 and g(x)=x-2, what is the domain of f(x)/g(x)?

OpenStudy (fortytherapper):

Domain is all of the x values except where x = 0, right? First solve for x x-2=0 After that, put the domain in interval notation (-Infinity, (Your answer for x)U((Your answer for x), infinity)

OpenStudy (whpalmer4):

Domain is all of the allowable values for the independent variable. Our function is \[\frac{f(x)}{g(x)} = \frac{(x+1)^{-1}}{x-2} = \frac{1}{(x+1)(x-2)}\] What are the values that \(x\) is not permitted to take?

OpenStudy (anonymous):

x+1=0 so x=-1 x-2=0 so x=2 Your domain can be any real value except -1 and 2